Crew 201 Astronomy Report 27Dec2018

[title Astronomy Report – December 27th]

[category astronomy-report]

Name: Federico Arturo Martinez Espinoza

Crew: Crew 201 MEx-1

Date: Dec 27 2018


Robotic Telescope Requested MDRS-WF

Objects Viewed: Messier 101

Problems Encountered: The filters used were RGB and H Alpha, the image taken with H alpha were not great and the final result after photoshop was not as expected, so I decided to remove the h alpha filter.

Greenhab Report – December 27th

Greenhab Officer:
Walter Calles

Environmental Control:

40% Shade Cloth on

Average Temperatures:
Low: 24°C
High: 33°C

Hours of Supplemental Light:

Daily Water Usage of Crops:
20.5 gallons

Water in Blue Tank (lbs):

Times of Watering for Crops:

Changes to crops:
Nothing significant

Re-planted half bed of Salad Savory Mix. Planted more cilantro and basil on blue flat. Planted more chives on the pot as well.
Added 2 new pots of onions and one new pot of spinach.
Re-arranged all the flats and pots on the greenhab for a better distribution.
Brought cucumbers to the front part of the greenhab and added some fertilizer. Hopefully, will see some progress in the next days.

Experiment Results:
Date: SOL7 (SOL12 overall)
Control Pot: 3 sprouts in good shape
Soil Sample Type 1
@10%: OK, 3 sprouts
@20%: OK, 3 sprouts
@30: OK, 2 sprouts
@40%: No progress
@50%: No progress

Soil Sample Type 2
@10%:OK, 2 sprouts
@20%:OK, 3 sprouts

Soil Sample Type 3
@10%:OK, 1 sprout
@20%:OK, 1 sprout


Support/Supplies Needed:
More sticky papers.

Operations Report – December 27th

SOL: 12

Name of person filing report: César Augusto Serrano Baza

Non-nominal systems: Generator Autostart

Notes on non-nominal systems: None
Generator (hours run): 16:55 yesterday until around 08:20. Today started at 17:11

Solar— SOC 78%
Diesel Reading – Around 74%
Propane Reading – 69%
Ethanol Free Gasoline – N/A gallons.

Water (Axillary tank) – 0 gallons
Water (static tank) – Medium to low level, around 100-gallons (See general notes and comments).
Axillary to Static tank transfer– no
Gallons transferred: 0
Water in GreenHab – Medium level (37%)
Water (loft) – Static to Loft Pump used – yes
Water Meter: 01395509

Toilet tank emptied: NO

Deimos rover used: no, still not functional
Hours: 0
Beginning charge: –
Ending charge: –
Currently charging: no

Sojourner rover used: ASSIGNED TO DIRECTOR
Beginning charge:
Ending charge:
Currently charging X :

Spirit rover used: NO
Hours: 63
Beginning charge: 100%
Ending charge: 100%
Currently charging: YES

Opportunity rover used: NO
Hours: 44.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: YES

Curiosity rover used: NO
Hours: 61.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: YES

Notes on rovers: Nothing to report.
# Hours the ATVs were Used today: Nothing to report.
Notes on ATVs : Nothing to report.

HabCar used and why, where? Not used.
CrewCar used and why, where? Not used.

General notes and comments: We had difficulties by pumping the water up from the static tank to the loft because the pipes were frozen. The director and us figured out that one end plug of the extention cord fo the pipe heater was not working. We changed the end plug and we were able to pump the water at 15:00 hrs.

Summary of internet: Nothing to report.

Summary of suits and radios: Nothing to report.

Summary of GreenHab operations: – Nothing to report.

Summary of ScienceDome operations: – Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Musk Solar observatory – Nominal. RObotic Observatory – Nominal.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: Nothing to report.

Thank you!!!

Sol Summary – December 27th

Crew 201 Sol Summary Report 27-Dec-2018

Sol: 12

Summary Title: This is about to come to its end.

Author’s name: Carlos Mariscal, XO.

Mission Status: Experiments are having some slow progress. Unfortunately, out VSAT, which is necessary to do one of the experiments, will not make it to the MDRS on time, so that experiment will not be performed. Depending on the progress of another of the experiments, we could end simulation tomorrow during the day or evening.

Sol Activity Summary: We:

– Continued activities of the artificial vision experiment.

– Printed several pieces for the rover.

– Had a lot of work at the GreenHab; we planted some seeds and organized some stuff.

– Recorded videos for media.

– Had a great time with some pictures in the snow (before it got kind of bad).

Look Ahead Plan: We hope to have our last EVA tomorrow to test mechanical components of the rover. Besides that, we will finish some experiments and start with the cleaning of the station for the next crew.

Anomalies in work: The plumb and water tank deicing system didn’t work properly, so we could not use any water during the morning. Fortunately, the temperature raised enough to get water up just before the snow started.

Weather: Cold in the morning, then snow, snow, wind and snow.

Crew Physical Status: A crew member had headaches, nothing serious, an aspirin solved it.

EVA: No EVA today.

Reports to be file:

– Commander report

– Sol summary

– HSO report

– Operations report

– GreenHab report

– Astronomer report

– Journalist report

– EVA request

– Photos

Support Requested: No support requested.

Copyright © The Mars Society. All rights reserved. | Main Site